OpenStudy (anonymous):

Find the equation of the tangent to the graph at the indicated point. f(x) = x2 − 4; a = 5 y=? I had the answer of 25x -124, but that's wrong :( help please.

OpenStudy (anonymous):

If x = 5, then substiute to find y, you get y = 21 Slope = derivative of f(x) = f'(x) = 2x so the slope is 2x and at x = 5, the slope is 2(5) = 10 Equation of tangent line is y - 21 = 10(x - 5) or, y - 21 = 10x - 50 or y = 10x - 29.